eeeeeeeeeeeeeeee:

What is the value of the expression shown below? 8 + (7 + 1) 2 ÷ 4 ⋅ 1 over 2 whole to the power of 4 (5 points) 7 9 21 24

sak123:

what do you think it is?

eeeeeeeeeeeeeeee:

9?

eeeeeeeeeeeeeeee:

I think...

eeeeeeeeeeeeeeee:

I don't really know.. but i think its 9

jhonyy9:

step by step please

jhonyy9:

@Laylalyssa

Xurlame:

Ok so, use PEMDAS to help you out

eeeeeeeeeeeeeeee:

ok

Xurlame:

If you go step by step using that you'll be able to find your answer

Xurlame:

Then tell us the answer you got and we'll help you out from there.

eeeeeeeeeeeeeeee:

ok i got 9

eeeeeeeeeeeeeeee:

as my final answer

eeeeeeeeeeeeeeee:

is it correct?

xxDeppressionxx:

What about 24 for the answer

eeeeeeeeeeeeeeee:

??? why? I got 9 ? How is it 24?

eeeeeeeeeeeeeeee:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxDeppressionxx What about 24 for the answer \(\color{#0cbb34}{\text{End of Quote}}\)

xxDeppressionxx:

No your right

xxDeppressionxx:

It is 9

xxDeppressionxx:

Add: 7 + 1 = 8 Exponentiation: the result of step No. 1 ^ 2 = 8 ^ 2 = 64 Divide: the result of step No. 2 / 4 = 64 / 4 = 16 Exponentiation: 1/ 2 ^ 4 = 14/ 24 = 1/ 16 In words - one half to four = one sixteenth. Multiple: the result of step No. 3 * the result of step No. 4 = 16 * 1/ 16 = 16 · 1/ 1 · 16 = 16/ 16 = 1 · 16/ 1 · 16 = 1 Multiply both numerators and denominators. Result fraction keep to lowest possible denominator GCD(16, 16) = 16. In the next intermediate step , cancel by a common factor of 16 gives 1/ 1 . In words - sixteen multiplied by one sixteenth = one. Add: 8 + the result of step No. 5 = 8 + 1 = 9